# Guided Diffusion Sampling on Function Spaces with Applications to PDEs

### v1

@@ 3.4 Approximation of Conditional Expectation via Tweedie’s Formula
Tweedie’s formula [53] plays a fundamental role in diffusion models by providing a link between score functions and conditional expectations. In infinite-dimensional spaces, we can no longer rely on probability density functions or standard gradients as in the finite-dimensional case. The notion of conditional expectation must be handled through measure-theoretic tools, and the score function needs to be redefined using the Fréchet derivative and the Riesz representation theorem.
Informally, our generalization (Theorem B.6) shows that under proper regularity assumptions, the conditional expectation of the initial state equals the score operator scaled by the noise variance:
$\mathbb{E}[{\boldsymbol{a}}_{0}\mid{\boldsymbol{a}}_{t}]=\sigma_{t}^{2}\mathbf{C}_{\gamma}\nabla_{{\boldsymbol{a}}_{t}}\log\frac{\,\mathrm{d}\nu_{t}}{\,\mathrm{d}\gamma_{t}}({\boldsymbol{a}}_{t}).$ (12)
This preserves the essential connection between scores and expectations from the finite-dimensional case. The full statement and proof are provided in Appendix B.
@@ 3.5 FunDPS: Function-Space Diffusion Posterior Sampling
With the extended Tweedie’s formula, we now have a closed form representation of $\hat{\boldsymbol{a}}_{0}({\boldsymbol{a}}_{t})=\mathbb{E}[{\boldsymbol{a}}_{0}\mid{\boldsymbol{a}}_{t}]$ given by Eq. 12. It becomes differentiable when we use a neural operator to simulate the score function, enabling us to compute the likelihood function given in Eq. 11.
In practice, our observations ${\boldsymbol{u}}$ will be finite dimensional. We also assume that the covariance $C_{\eta}$ of the observation noise is approximately uncorrelated, such that $C_{\eta}^{-1}$ can be well-approximated by a scaled version $C_{\eta}^{-1}\approx cI$ of the identity. Plugging this and Eq. 11 into Eq. 6 with the trained score operator ${\boldsymbol{D}}_{\theta}$, we can now formalize the conditional score operator, acting as the guidance for posterior sampling
$\nabla_{{\boldsymbol{a}}_{t}}\log\frac{\,\mathrm{d}\nu_{t}^{{\boldsymbol{u}}}}{\,\mathrm{d}\gamma_{t}}({\boldsymbol{a}}_{t})=\nabla_{{\boldsymbol{a}}_{t}}\log\frac{\,\mathrm{d}\nu_{t}}{\,\mathrm{d}\gamma_{t}}({\boldsymbol{a}}_{t})+\nabla_{{\boldsymbol{a}}_{t}}\tilde{\Phi}_{t}({\boldsymbol{a}}_{t},{\boldsymbol{u}})\approx{\boldsymbol{D}}_{\theta}({\boldsymbol{a}}_{t},t)-\tfrac{c^{2}}{2}\nabla_{{\boldsymbol{a}}_{t}}\lVert{\boldsymbol{u}}-{\boldsymbol{A}}(\hat{{\boldsymbol{a}}}_{0})\rVert_{\mathcal{U}}^{2}.$ (13)
{algorithm}
[…]
Figure 3: FunDPS Sampler
[…]
@@
Algorithm 3 depicts this reverse diffusion process, where we iteratively update the samples ${\boldsymbol{a}}_{i}$ using the consistency between the given measurement ${\boldsymbol{u}}$ and the one obtained from the denoiser, i.e., ${\boldsymbol{A}}({\boldsymbol{D}}_{\theta}({\boldsymbol{a}}_{i},t_{i}))$. Specifically, we propose the update rule
@@
Beyond reducing steps, inference time per step can be accelerated by initially processing at a lower resolution and progressively increasing it. This aligns with the concept of diffusion models generating autoregressively in resolution or frequency space [57]. We propose ReNoise, a bi-level multi-resolution inference method. Initially, samples are processed at a lower resolution. Subsequently, they are upscaled, and, inspired by ReSample[58], additional noise is introduced to mitigate artifacts and noise level mismatch arising from upscaling, followed by a few denoising steps at the target resolution. This enables us to achieve equivalent accuracy in half the computing time. We illustrate the pipeline in Figure 4b with implementation details in Section G.4.
@@ 4 Experiments
We will now showcase the efficiency, robustness, and multi-resolution generalizability of our proposed FunDPS through various tasks. Additional and ablation experiments are in Appendix G.
@@ 4.1 Experimental Setup
We validate our approach by solving both forward and inverse problems on five different PDE problems. The problems span a range of difficulties, characterized by complex input distributions arising from the multi-scale features of GRFs, varying boundary conditions (Dirichlet and periodic), and nonlinear patterns inherent in the Navier-Stokes equations. We provide a detailed description of each PDE in Appendix D. The objective is to solve forward and inverse problems in sparse sensor settings. For forward tasks, we only observe 3% of the initial condition before reconstructing the entire solution. Conversely, for inverse tasks, we only observe 3% of the final state before reconstructing unknown initial or coefficient fields.
@@
We base our code on EDM-FS [59]. In particular, we adopt a U-shaped neural operator architecture [52] as the denoiser ${\boldsymbol{D}}_{\theta}$ and modify the noising process according to the discussion above. Our denoiser has 54M parameters, which is similar to DiffusionPDE’s network size. The implementation is detailed in Appendix E. The guidance weight $\zeta$ in Eq. 14 is tuned on a small validation set, provided in Appendix F.
@@
Baseline comparisons include well-known deterministic PDE solvers such as FNO [21], PINO [60], DeepONet [22] and PINN [18], as well as the state-of-the-art diffusion-based approach, DiffusionPDE [17]. We had correspondence with DiffusionPDE’s authors about its reproducibility issues. Please refer to the Appendix I for details.
@@ 4.2 Results
We evaluate our method on both forward and inverse problems across five PDE tasks with sparse observations. Results are shown in Table 1. Even with severe occlusions, where only $3\%$ of function points are visible, we are able to reconstruct both states effectively. Our approach always achieves the best results among all the baselines with the lowest error rates, surpassing all baselines with 32% higher accuracy with one-fourth steps. We provide qualitative results in Figure 1 and Appendix H. We also compared all methods on classical fully-observed forward and inverse problems, where FunDPS also showed superior performance as in Table 6. Further, we test our framework with diverse guidance methods to show its adaptability in Section G.1 (Figure 6 and Table 5). We also conducted an ablation study on the effectiveness of our design choices in Table 9. Additionally, we provide results by using FDM as forward operator and compare with joint learning in Appendix C.
@@
We compare the inference time of our method with the diffusion-based solver, DiffusionPDE, as shown in Figure 4. With a 200-step discretization of the reverse-time SDE, our method achieved superior accuracy with only one-tenth of the integration steps compared to DiffusionPDE. When we increase the number of steps, FunDPS further reduces the error, which suggests that the scaling law of inference time may hold for solving PDE problems with guided diffusion. This highlights the effectiveness of our function space formulation and guidance mechanism.
@@
Multi-resolution training and inference
For our main results, we employed a two-phase training strategy: training primarily on low-resolution data before switching to high resolution in the final epochs. As shown in Table 7, this approach achieves comparable accuracy to models trained solely on high-resolution grids, while requiring only $25\%$ of the GPU hours. We also assessed our multi-resolution inference technique, ReNoise. With $80\%$ of steps performed at low resolution, ReNoise sustains similar accuracy (see Figure 7), yielding a further 2x speed improvement to 7.5s/sample–25 times faster than DiffusionPDE.
@@ A.1 Forward Diffusion with Gaussian Random Fields
Noising process: Given a sample $a\sim\mu$, we perturb it by adding an independent Gaussian random function (drawn from the GRF prior) with appropriate variance. In other words, for a noise scale $\sigma\geq 0$, we define the noisy function at level $\sigma$ as
@@ A.2 Score Function and Denoising Objective
Denoising score matching: In practice, $s(a,\sigma)$ is unknown because the true data distribution $\mu$ (and hence $\mu_{\sigma}$) is unknown. Instead of trying to directly estimate the score, we train a denoising model $D_{\theta}(a_{\sigma},\sigma)$ to recover the underlying clean function $a$ from a noisy sample $a_{\sigma}$. This is called denoising score matching, which is particularly convenient in function spaces. Following the EDM [16] training strategy, we sample pairs of clean and noisy functions and train $D_{\theta}$ to predict the clean input. In particular, given a sample $a\sim\mu$ and its noisy version $y=a+\eta$ with $\eta\sim\mathcal{N}(0,\sigma^{2}\mathbf{C})$, we train $D_{\theta}$ to output $a$ (the ground truth) when given $(y,\sigma)$ as input. The training objective, averaged over the data distribution and noise, is formulated as a weighted mean squared error:
@@
Thus, by training the model to minimize $L(\theta)$ across many noise levels, we are effectively teaching it to approximate the score operator in $H$. We implement the above training with a discrete noise level schedule and random sampling of noise intensities, similar to the EDM methodology (Algorithm A.2). We randomly sample $\sigma$ noise levels log-uniformly over $[\sigma_{\min},\,\sigma_{\max}]$ as in Karras et al. [16]. As a result, $D_{\theta}$ becomes a function-space denoiser that can gradually refine a noisy input at any noise scale within the range.
@@ A.3 Reverse Diffusion and Sampling
Once the denoising model (score model) is trained, we can generate new function samples from the learned distribution by running the diffusion process in reverse – starting from noise and iteratively removing noise. The key idea is to start with an initial random field drawn from the prior and then repeatedly apply the denoiser $D_{\theta}$ while decreasing $\sigma$ in stages. Here we adopt a discrete reverse diffusion approach aligned with EDM’s deterministic solver. First, we choose a high noise level $\sigma_{\max}$ (e.g. the upper bound used in training) and sample an initial function $a_{N}\sim\mathcal{N}(0,;\sigma_{\max}^{2},\mathbf{C})$, i.e. a pure noise sample in $H$ drawn from the GRF prior. Then we define decreasing sequence of noise levels $\sigma_{\max}=\sigma_{N}>\sigma_{N-1}>\cdots>\sigma_{1}>\sigma_{0}\approx 0$, where the levels are spaced polynomially so that adjacent levels have small differences in terms of signal-to-noise ratio as in EDM [16]. The noise scheduler spans from the highest noise to zero ($\sigma_{\max}$ corresponds to the prior and $\sigma_{0}=0$ corresponds to a clean sample). At each step $i=N,N-1,\dots,1$, given the current noisy sample $a_{i}$ at noise level $\sigma_{i}$, we apply the denoiser to get $\hat{a}=D_{\theta}(a_{i},\sigma_{i})$, the model’s estimate of the clean underlying function. It is followed by adding the $\sigma_{i-1}$ noise level to reach the next step’s sample $a_{i-1}$ along with higher-order updates. After iterating down to the final level $\sigma_{0}\to 0$, we obtain $a_{0}$, which is an approximate sample from the original data distribution $\mu$ (since no noise remains). We can then use $a_{0}$ as a newly generated function sample drawn from the learned generative model. We implement our framework using a deterministic sampler based on Euler’s $2^{\text{nd}}$ order method. Algorithm 3 without the FunDPS guidance corresponds to this unconditional sampling procedure.
Appendix B Measure-Theoretic Tweedie’s Formula
@@
B.1 Banach Space Gaussian Measures
@@
Although $H(\gamma)\subset B^{**}$ in general, because $B$ is a separable Banach space we may take $H(\gamma)\subset B$. That is, we write $h=C_{\gamma}(f)$ for some $h\in B$ and $f\in B^{*}_{\gamma}$ if $f(g)=C_{\gamma}(f)(g)$ for all $g\in B^{*}$. Since every $h,k\in H(\gamma)$ are of the form $h=C_{\gamma}(\hat{h})$, $k=C_{\gamma}(\hat{k})$ for some $\hat{h},\hat{k}\in B^{*}_{\gamma}$, the space $H(\gamma)$ has an induced inner product
@@
Theorem B.1 (Cameron-Martin).
@@
$\frac{\,\mathrm{d}\gamma^{h}}{\,\mathrm{d}\gamma}(x)=\exp\left(C_{\gamma}^{-1}(h)(x)-\tfrac{1}{2}|h|^{2}_{H(\gamma)}\right).$
B.2 Tweedie’s Formula
@@
We begin by showing that $\nu\sim\gamma$ are equivalent when $\mu$ is supported on the Cameron-Martin space $H(\gamma)$. This is a generalization of [14, Theorem 1], who show an analogous claim for separable Hilbert spaces. The proof is essentially the same in both cases, but we include it here for the sake of completeness.
Proposition B.2.
@@
Let $E\in\mathcal{B}(B)$ be an arbitrary Borel set. Suppose that $\gamma(E)=0$. By Theorem B.1, $\gamma^{x}\sim\gamma$ for $\mu$-almost every $x$, and hence $\gamma^{x}(E)=0$ for $\mu$-almost every $x$. It follows that
@@
Conversely, suppose $\nu(E)=0$. It follows that $\gamma^{x}(E)=0$ for $\mu$-almost every $x$. Since $\mu(H(\gamma))=1$, Theorem B.1 shows the measures $\gamma^{x}$ and $\gamma$ are almost surely equivalent and thus $\gamma(E)=0$. ∎
@@
That is, $D_{H(\gamma)}\phi=D_{H(\gamma)}\log\frac{\,\mathrm{d}\nu}{\,\mathrm{d}\gamma}$ is the logarithmic derivative of the density of $\nu$ along $H(\gamma)$. The value $[D_{H(\gamma)}\phi](x)(h)$ is the derivative of $\phi$ at $x$ in the direction $h\in H(\gamma)$. We refer to [14] for a further discussion of this notion of a score and the differentiability assumption.
@@
where the last equality follows by Tonelli’s theorem and the fact that the densities are nonnegative. By Proposition B.2, $\nu\ll\gamma$ and so by the Radon-Nikodym theorem the density $\,\mathrm{d}\nu/\,\mathrm{d}\gamma$ is uniquely defined up to a set of $\gamma$-measure zero. Thus,
@@
We now proceed to calculate the Cameron-Martin space gradient of the score. In particular, we show that it is equal to the same expression we obtained in Proposition B.4. This requires an assumption on the measure $\mu$ to ensure that the derivatives of $\,\mathrm{d}\gamma^{x}/\,\mathrm{d}\gamma$ are bounded by an integrable function in order to justify a derivative-integral exchange. Using the continuity of $C_{\gamma}$, the condition in Eq. 43 can be relaxed to finding an integrable $\psi\in L^{1}(\gamma)$ such that for all $y\in B$ and $\mu$-almost every $x$,
@@
where $J:B^{*}_{\gamma}\to H^{*}$ is the isomorphism defined in Eq. 24. Note this isomorphism is required as $C_{\gamma}^{-1}(x)\in B^{*}\gamma$ is an element of the RKHS $B^{*}_{\gamma}$, whereas the Fréchet derivative is an element of $H(\gamma)^{*}$. Using Lemma B.3 and the assumption that there exists a dominating function $\psi\in L^{1}(\mu)$, we may use the Leibniz integral rule to calculate that
@@
Combined with Eq. 46, this yields the claim. ∎
Combining Proposition B.4 and Proposition B.5 yields our Banach space generalization of Tweedie’s formula below.
Theorem B.6.
B.3 Special Case: Euclidean Setting

### v2

@@ 3.4 Approximation of Conditional Expectation via Tweedie’s Formula
Tweedie’s formula [54] plays a fundamental role in diffusion solvers by providing a link between score functions and conditional expectations. It is a cornerstone in many guided sampling methods, including MCG [38], DPS [40], and PSLD [55]. In our framework, the approximated likelihood also relies on this efficient estimation of conditional expectations $\mathbb{E}[{\boldsymbol{a}}_{0}\mid{\boldsymbol{a}}_{t}]$, but in function spaces.
Generalization of Tweedie’s formula to function spaces requires careful treatment of several concepts. In infinite-dimensional spaces, we can no longer rely on probability density functions or standard gradients as in the finite-dimensional case. The notion of conditional expectation must be handled through measure-theoretic tools, and the score function needs to be redefined using the Fréchet derivative and the Riesz representation theorem. Here we present our extension below.
Theorem 3.1 (Tweedie’s formula in infinite-dimensional Banach spaces).
Proof.
See Appendix˜C. ∎
@@ 3.5 FunDPS: Function-Space Diffusion Posterior Sampling
With the extended Tweedie’s formula, we now have a closed form representation of $\hat{\boldsymbol{a}}_{0}({\boldsymbol{a}}_{t})=\mathbb{E}[{\boldsymbol{a}}_{0}\mid{\boldsymbol{a}}_{t}]$ given by Theorem˜3.1. It becomes differentiable when we use a neural operator to simulate the score function, enabling us to compute the likelihood function given in Eq.˜11.
In practice, our observations $\boldsymbol{u}$} will be finite dimensional. We also assume that the covariance $C_{\eta}$ of the observation noise is approximately uncorrelated, such that $C_{\eta}^{-1}$ can be well-approximated by a scaled version $C_{\eta}^{-1}\approx cI$ of the identity. Plugging this and Eq.˜11 into Eq.˜6 with the trained score operator ${\boldsymbol{D}}_{\theta}$, we can now formalize the conditional score operator, acting as the guidance for posterior sampling
$\nabla_{{\boldsymbol{a}}_{t}}\log\frac{\,\mathrm{d}\nu_{t}^{{\boldsymbol{u}}}}{\,\mathrm{d}\gamma_{t}}({\boldsymbol{a}}_{t})=\nabla_{{\boldsymbol{a}}_{t}}\log\frac{\,\mathrm{d}\nu_{t}}{\,\mathrm{d}\gamma_{t}}({\boldsymbol{a}}_{t})+\nabla_{{\boldsymbol{a}}_{t}}\tilde{\Phi}_{t}({\boldsymbol{a}}_{t},{\boldsymbol{u}})\approx{\boldsymbol{D}}_{\theta}({\boldsymbol{a}}_{t},t)-\tfrac{c}{2}\nabla_{{\boldsymbol{a}}_{t}}\lVert{\boldsymbol{u}}-{\boldsymbol{A}}(\hat{{\boldsymbol{a}}}_{0})\rVert_{\mathcal{U}}^{2}.$ (13)
@@
Algorithm B depicts this reverse diffusion process, where we iteratively update the samples ${\boldsymbol{a}}_{i}$ using the consistency between the given measurement $\boldsymbol{u}$} and the one obtained from the denoiser, i.e., ${\boldsymbol{A}}({\boldsymbol{D}}_{\theta}({\boldsymbol{a}}_{i},t_{i}))$. Specifically, we propose the update rule
@@
Beyond reducing steps, inference time per step can be accelerated by initially processing at a lower resolution and progressively increasing it. This aligns with the concept of diffusion models generating autoregressively in resolution or frequency space [59]. We propose ReNoise, a bi-level multi-resolution inference method. Initially, samples are processed at a lower resolution. Subsequently, they are upscaled, and inspired by ReSample[60], additional noise is introduced to mitigate artifacts and noise level mismatch arising from upscaling, followed by a few denoising steps at the target resolution. This enables us to achieve equivalent accuracy in half the computing time. We illustrate the pipeline in Figure˜3b with implementation details in Section˜H.4.
@@ 4 Experiments
We will now showcase the efficiency, robustness, and multi-resolution generalizability of our proposed FunDPS through various tasks. Additional and ablation experiments can be found in Appendix˜H.
@@ 4.1 Experimental Setup
We validate our approach by solving both forward and inverse problems on five different PDE problems. The problems span a range of difficulties, characterized by complex input distributions arising from the multi-scale features of GRFs, varying boundary conditions (Dirichlet and periodic), and nonlinear patterns inherent in the Navier-Stokes equations. We provide a detailed description of each PDE in Appendix˜E. The objective is to solve forward and inverse problems in sparse sensor settings. For forward tasks, we only observe 3% of the initial condition before reconstructing the entire solution. Conversely, for inverse tasks, we only observe 3% of the final state before reconstructing unknown initial or coefficient fields.
@@
We base our code on EDM-FS [61]. In particular, we adopt a U-shaped neural operator architecture [52] as the denoiser ${\boldsymbol{D}}_{\theta}$ and modify the noising process according to the discussion above. Our denoiser has 54M parameters, which is similar to DiffusionPDE’s network size. The implementation is detailed in Appendix˜F. The guidance weight $\zeta$ in Eq.˜14 is tuned on a small validation set, provided in Appendix˜G.
@@
Baseline comparisons include well-known deterministic PDE solvers such as FNO [21], PINO [62], DeepONet [22] and PINN [18], as well as the state-of-the-art diffusion-based approach, DiffusionPDE [17]. We had correspondence with DiffusionPDE’s authors about its reproducibility issues. We use our reproduced results for comparison. Please refer to the Appendix˜I for details.
@@ 4.2 Results
We evaluate our method on both forward and inverse problems across five PDE tasks with sparse observations. Results are shown in Table˜1. Even with severe occlusions, where only $3\%$ of function points are visible, we are able to reconstruct both states effectively. Our approach always achieves the best results among all the baselines with the lowest error rates, surpassing all baselines with 32% higher accuracy with one-fourth steps. We provide qualitative results in Figure˜1 and Appendix˜K. We also compared all methods on classical fully-observed forward and inverse problems, where FunDPS also showed superior performance as in Table˜6. Further, we test our framework with diverse guidance methods to show its adaptability (Figure˜6 and Table˜5). We also conducted an ablation study on the effectiveness of our design choices in Table˜12. Additionally, we provide results by using FDM as forward operator and compare with joint learning in Appendix˜D.
@@
We compare the inference time of our method with the diffusion-based solver, DiffusionPDE, as shown in Figure˜3. With a 200-step discretization of the reverse-time SDE, our method achieved superior accuracy with only one-tenth of the integration steps compared to DiffPDE. When we increase the number of steps, FunDPS further reduces the error, which suggests that the scaling law of inference time may hold for solving PDE problems with guided diffusion. This highlights the effectiveness of our function space formulation and guidance mechanism.
@@
Multi-resolution training
For our main results, we employed a two-phase training strategy: training primarily on low-resolution data before switching to high resolution in the final epochs. As shown in Table˜7, this approach achieves comparable accuracy to models trained solely on high-resolution grids, while requiring only $25\%$ of the GPU hours.
Multi-resolution inference
Thanks to its multi-resolution nature, neural operators can be trained on multiple resolutions and applied to data of different resolutions. We can reduce significant inference time by performing most sampling steps at low resolution and upsampling only near the end to finalize high-frequency details. We found that upsampling in the middle of the diffusion process worked poorly, which we attribute to the difficulty of preserving GRF properties during upscaling. Hence, we propose a bi-level sampling process, ReNoise, that mitigates the upsampling artifacts by adding noise for improved correction potential. The implementation details are given in Section˜H.4.
We provide the multi-resolution inference results in Figure˜4. With $80\%$ of steps performed at low resolution, ReNoise sustains similar accuracy, yielding a further 2x speed improvement to 7.5s/sample–25 times faster than DiffusionPDE.
@@ A.1 Forward Diffusion with Gaussian Random Fields
Noising process
Given a sample $a\sim\mu$, we perturb it by adding an independent Gaussian random function (drawn from the GRF prior) with appropriate variance. In other words, for a noise scale $\sigma\geq 0$, we define the noisy function at level $\sigma$ as
@@ A.2 Score Function and Denoising Objective
Denoising score matching
In practice, $s(a,\sigma)$ is unknown because the true data distribution $\mu$ (and hence $\mu_{\sigma}$) is unknown. Instead of trying to directly estimate the score, we train a denoising model $D_{\theta}(a_{\sigma},\sigma)$ to recover the underlying clean function $a$ from a noisy sample $a_{\sigma}$. This is called denoising score matching, which is particularly convenient in function spaces. Following the EDM [16] training strategy, we sample pairs of clean and noisy functions and train $D_{\theta}$ to predict the clean input. In particular, given a sample $a\sim\mu$ and its noisy version $y=a+\eta$ with $\eta\sim\mathcal{N}(0,\sigma^{2}\mathbf{C})$, we train $D_{\theta}$ to output $a$ (the ground truth) when given $(y,\sigma)$ as input. The training objective, averaged over the data distribution and noise, is formulated as a weighted mean squared error:
@@
Thus, by training the model to minimize $L(\theta)$ across many noise levels, we are effectively teaching it to approximate the score operator in $H$. We implement the above training with a discrete noise level schedule and random sampling of noise intensities, similar to the EDM methodology (Algorithm A.2). We randomly sample $\sigma$ noise levels log-uniformly over $\sigma_{\min},\,\sigma_{\max}]$ as in Karras et al. [16]. As a result, $D_{\theta}$ becomes a function-space denoiser that can gradually refine a noisy input at any noise scale within the range.
@@ A.3 Reverse Diffusion and Sampling
Once the denoising model (score model) is trained, we can generate new function samples from the learned distribution by running the diffusion process in reverse – starting from noise and iteratively removing noise. The key idea is to start with an initial random field drawn from the prior and then repeatedly apply the denoiser $D_{\theta}$ while decreasing $\sigma$ in stages. Here we adopt a discrete reverse diffusion approach aligned with EDM’s deterministic solver. First, we choose a high noise level $\sigma_{\max}$ (e.g. the upper bound used in training) and sample an initial function $a_{N}\sim\mathcal{N}(0,;\sigma_{\max}^{2},\mathbf{C})$, i.e. a pure noise sample in $H$ drawn from the GRF prior. Then we define decreasing sequence of noise levels $\sigma_{\max}=\sigma_{N}>\sigma_{N-1}>\cdots>\sigma_{1}>\sigma_{0}\approx 0$, where the levels are spaced polynomially so that adjacent levels have small differences in terms of signal-to-noise ratio as in EDM [16]. The noise scheduler spans from the highest noise to zero ($\sigma_{\max}$ corresponds to the prior and $\sigma_{0}=0$ corresponds to a clean sample). At each step $i=N,N-1,\dots,1$, given the current noisy sample $a_{i}$ at noise level $\sigma_{i}$, we apply the denoiser to get $\hat{a}=D_{\theta}(a_{i},\sigma_{i})$, the model’s estimate of the clean underlying function. It is followed by adding the $\sigma_{i-1}$ noise level to reach the next step’s sample $a_{i-1}$ along with higher-order updates. After iterating down to the final level $\sigma_{0}\to 0$, we obtain $a_{0}$, which is an approximate sample from the original data distribution $\mu$ (since no noise remains). We can then use $a_{0}$ as a newly generated function sample drawn from the learned generative model. We implement our framework using a deterministic sampler based on Euler’s $2^{\text{nd}}$ order method. Algorithm B without the FunDPS guidance corresponds to this unconditional sampling procedure.
Appendix B Pseudocode for FunDPS
{algorithm}
[…]
Appendix C Measure-Theoretic Tweedie’s Formula
@@
C.1 Banach Space Gaussian Measures
@@
[…]
@@
Theorem C.1 (Cameron-Martin).
@@
[…]
C.2 Tweedie’s Formula
@@
We begin by showing that $\nu\sim\gamma$ are equivalent when $\mu$ is supported on the CM space $H(\gamma)$. This is a generalization of [14, Theorem 1], who show an analogous claim for separable Hilbert spaces. The proof is essentially the same in both cases, but we include it here for the sake of completeness.
Proposition C.2.
@@
Let $E\in\mathcal{B}(B)$ be an arbitrary Borel set. Suppose that $\gamma(E)=0$. By Theorem C.1, $\gamma^{x}\sim\gamma$ for $\mu$-almost every $x$, and hence $\gamma^{x}(E)=0$ for $\mu$-almost every $x$. It follows that
@@
Conversely, suppose $\nu(E)=0$. It follows that $\gamma^{x}(E)=0$ for $\mu$-almost every $x$. Since $\mu(H(\gamma))=1$, Theorem C.1 shows the measures $\gamma^{x}$ and $\gamma$ are almost surely equivalent and thus $\gamma(E)=0$. ∎
@@
That is, $D_{H(\gamma)}\phi=D_{H(\gamma)}\log\frac{\,\mathrm{d}\nu}{\,\mathrm{d}\gamma}$ is the logarithmic derivative of the density of $\nu$ along $H(\gamma)$. The value $D_{H(\gamma)}\phi](x)(h)$ is the derivative of $\phi$ at $x$ in the direction $h\in H(\gamma)$. We refer to [14] for a further discussion of this notion of a score and the differentiability assumption.
@@
where the last equality follows by Tonelli’s theorem and the fact that the densities are nonnegative. By Proposition C.2, $\nu\ll\gamma$ and so by the Radon-Nikodym theorem the density $\mathrm{d}\nu/\,\mathrm{d}\gamma$ is uniquely defined up to a set of $\gamma$-measure zero. Thus,
@@
We now proceed to calculate the Cameron-Martin space gradient of the score. In particular, we show that it is equal to the same expression we obtained in Proposition C.4. This requires an assumption on the measure $\mu$ to ensure that the derivatives of $\mathrm{d}\gamma^{x}/\,\mathrm{d}\gamma$ are bounded by an integrable function in order to justify a derivative-integral exchange. Using the continuity of $C_{\gamma}$, the condition in Eq.˜43 can be relaxed to finding an integrable $\psi\in L^{1}(\gamma)$ such that for all $y\in B$ and $\mu$-almost every $x$,
@@
where $J:B^{*}_{\gamma}\to H^{*}$ is the isomorphism defined in Eq.˜24. Note this isomorphism is required as $C_{\gamma}^{-1}(x)\in B^{*}\gamma$ is an element of the RKHS $B^{*}_{\gamma}$, whereas the Fréchet derivative is an element of $H(\gamma)^{*}$. Using Lemma C.3 and the assumption that there exists a dominating function $\psi\in L^{1}(\mu)$, we may use the Leibniz integral rule to calculate that
@@
Combined with Eq.˜46, this yields the claim. ∎
Combining Proposition C.4 and Proposition C.5 yields the Banach space generalization of Tweedie’s formula, which concludes the proof for Theorem˜3.1.
C.3 Special Case: Euclidean Setting